# Evaluating LLMs on microservice-based applications: how complex is your specification?

### v1

@@ 5.2 Installing packages
We run the LLM generated code in a Docker container. The Dockerfile for that container contains a command to install dependencies in the requirements.txt file using pip. We generate the requirements.txt file by scanning the generated code for import statements [^3] In our prompt we asked the LLM to document which packages to install. But we found that it would make mistakes and it was a better solution to install packages based upon the import statements. . This was problematic as pip would fail when installing a package that comes with the standard python distribution (such as uuid or json). To get around this problem, we omitted packages included in the Python standard library.
In the restaurant application (Section 5.4), the generated code for the authentication MS would always import the jwt package, but it required the installation of pyJWT. (To its credit, the LLM would note this in a comment.) For this and other similar issues, our code substitutes the package name in the import statement with the correct package. Nonetheless, as we show below and documented by other research [24, 12, 9], choosing the right package and version was still problematic.
@@ 5.3 The Public Library Application
Our first MSBA is a public library application comprised of 4 MSs: Cardholders, Books, Borrows, and Logs. The Cardholders service was given in Figure 1. Books is responsible for managing the books in the library, Borrows is responsible for tracking books on loan, and Logs keeps a record of each book that is borrowed from the library. The specification of these MSes can be found on-line (see footnote 2).
We test this application using 3 different LLMs: GPT-3.5-16K, GPT-4o-mini, and GPT-o3-mini. For each LLM we performed two sets of experiments, one using a 0-shot prompt, and one using a $1$-shot prompt [5]. The $1$-shot prompt, given in Appendix C, provides example code using Flask and MongoDB for a small MS. For each LLM and each MS of the MSBA we repeat each experiment 15 times. For GPT-3.5 and GPT-4o-mini we perform 5 experiments for each temperature $0.0$, $0.3$, and $0.5$. GPT-o3-mini did not support different temperature settings when we ran our experiments and hence all experiments used the default temperature. Each experiment tests the MSBA with unit tests for each MS. There are 15 unit tests for Cardholders, 14 for Books and Borrows, and 7 for Logs. The latter has fewer tests because Logs only allows external actors to perform GET requests; they cannot create, delete or modify logs on their own. For each experiment, if the LLM generated code does not pass all the unit tests, we prompt the LLM to reflect on the nature of the errors and then ask it to regenerate code, as explained above. We only do this once (MaxGen = 2 in Figure 2). We refer to the initial generated code as V0 and the regenerated code as V1.
@@ 5.3.1 Results
Tables 1 and 2 shows the results of our experiments on GPT-3.5 with the 0-shot and $1$-shot prompts respectively. The $1^{st}$ column is the name of the MS. As not all code generated by the LLM was executable, the $2^{nd}$ column lists the fraction of the 15 experiments that generated executable V0 code permitting execution of the unit tests. The $3^{rd}$ column lists the number of experiments that generated executable V1 code. The $4^{th}$ and $5^{th}$ columns present the percent of unit tests that passed using V0 and V1 code respectively, averaged over all experiments that produced executable code444 We do not consider unexecutable generated code in this percentage. We wanted our results to focus on how often executable code was functionally correct. . The $6^{th}$ column is of the form $n_{1}$-$n_{2}/t$, where $n_{1}$ ($n_{2}$) is the highest number of unit tests passed in any experiment for V0 (V1) and $t$ is the total number of unit tests for that MS. The last column is the percent of experiments in which the number of unit tests passed by V1 was greater than the number of unit tests passed by V0, considering only those experiments in which V1 was executed.
Name # V0 # V1 # V0 % V1 Highest % V1
@@
Table 1: Library MSBA results using GPT-3.5 with 0-shot prompt
Name # V0 # V1 # V0 % V1 Highest % V1
@@
Table 2: Library MSBA results using GPT-3.5 with 1-shot prompt
@@
Table 3: Library MSBA results using GPT-o3-mini with 0-shot prompt
@@
Table 4: Library MSBA results using GPT-o3-mini with 1-shot prompt
We omit the results for GPT-4o-mini due to space considerations. In general it performs better than GPT-3.5 but not as good as GPT-o3-mini. See Appendix D.
@@ 5.3.2 Discussion
A third observation, based upon column 4 of Tables 2 and 4, is that providing an example (1-shot prompt) significantly improves performance of GPT-3.5 but mostly degrades performance with GPT-o3-mini. This is often the case with reasoning models like GPT-o3-mini [^5] For instance, the creators of DeepSeek -R1 report that “When evaluating DeepSeek-R1, we observe that it is sensitive to prompts. Few-shot prompting consistently degrades its performance. Therefore, we recommend users directly describe the problem and specify the output format using a zero-shot setting for optimal results.” [7] See also [27]. , and especially in our case, where performance is high without the example, and the example can cause overfitting.
@@
A general comment is warranted on these statistics; namely, not all unit tests are equal. Some unit tests may require complicated logic while others just check very simple REST patterns that do not even require knowledge of the MS. For instance, one of the unit tests checks that a POST request with a non-JSON payload fails. Hence while percentage of unit tests passed still gives us an overall idea on the strength of the LLM, we require a deeper analysis examining which tests pass or fail to obtain more insight (see Section 5.7).
With respect to temperature, while the $0.3$ setting often performs best for V0 and $0.5$ for V1, the differences in performance between different temperatures is usually less than $10\%$. See Appendix D for details.
@@ 5.4 The Restaurant Application
Given the poor results of GPT-3.5 on the Library app, we use only GPT-4o-mini and GPT-o3-mini for our experiments on this MSBA. For similar reasons, we only experiment with a 0-shot prompt. For GPT-4o-mini we performed 20 experiments for each MS. For the ratings MS, 8 of the experiments were at temperature $0.0$, 7 were at temperature $0.3$ and 5 were at temperature $0.5$. For all other microservices, there were 10 experiments at temperature $0.0$, and 5 experiments at each of the temperatures $0.3$ and $0.5$. For GPT-o3-mini we performed 15 experiments for each MS. As mentioned above GPT-o3-mini did not support different temperatures at the time of our experiments, so we used the default temperature.
Each experiment consists of 9, 15, 10 and 8 unit tests for the Authentication, Dishes, Profile, and Ratings MSes respectively. As above, for each experiment, if the LLM generated code does not pass all unit tests, we ask the LLM to reflect on the nature of the errors and then regenerate code.
@@ 5.6 Description Complexity
Before analyzing the results of these experiments, we first examine the complexity of the restaurant and library MSBAs, given in Table 7. The description size is for the 0-shot prompt. The 1-shot prompt adds $490$ words.
The Restaurant MSBA has a description size about .36.36 larger than the Library application, and has twice the number of dependencies, although the package complexity is the same. This at least partially explains the difference in LLM performance between the library and restaurant MSBAs.
Name Description # Dependencies Avg # Packages LLM assigned
size (words) per MS difficulty
Library 1399 3 4 3
Restaurant 1905 6 4 4
@@
The LLM-as-a-Judge score, provided by Claude Sonnet 4, confirms this conclusion. It assigns the library application a difficulty of 3 (out of 5) while the restaurant application received a complexity score of 4. It says that the library app has “moderate complexity” and that it is not overly complex because of the “Clear, well-defined requirements”, “Standard CRUD operations”, “Common technology stack”, “Simple data models”, and “Limited business logic”. On the other hand, factors that add complexity include “Inter-service communication”,“External API integration”, “Business rule enforcement”, “Date calculations”, “Error handling”, and “Distributed data consistency”.
For the restaurant MSBA, the LLM states the following “High Complexity Elements” contributing to the 4 rating: “Multi-service Architecture”, “Authentication & Authorization”, “External API Integration”, “Complex Business Logic”, “Database Operations”, “Error Handling”, and “Query Processing”. It did not give it a 5 rating because of “Well-defined specifications with clear requirements”, “Standard technologies (Python, MongoDB, REST APIs)”, “No real-time processing or complex distributed systems patterns” and “No advanced security beyond basic token auth”. In Appendix B we provide the LLM-as-a-Judge prompt and the complete comments provided by the LLM to explain its score.
@@ 5.7 Discussion
From Table 5 we see that GPT-4o-mini does extremely poorly on the restaurant application, and while GPT-o3-mini (Table 6) does better, it still never performs above $25\%$ passing unit tests with V0 code except for the Authentication MS, where it achieves $65.8\%$. We also note that neither V0 nor V1 code ever achieves a perfect score, with the exception of the Authentication service. The V1 code of that service, as seen in column $6$, sometimes passes all $9$ unit tests. This is true for 6 out of the 11 V1 testable implementations.
To get a deeper understanding of the cause of the errors, we examined the run-time error-logs, the GPT-o3-mini LLM-generated reflection results, as well as the generated code. Some of the persistent problems we found in the code are the following:
  • Incorrect package usage. For instance, even though we installed pyJWT (see Section 5.2 ), the authentication code would sometimes use this package incorrectly. This might be due to an incorrect version being installed. Along similar lines, Ratings would consistently import the flask_jwt_extended package. A typical error message was the following:
@@
    We tried multiple versions of the flask_jwt_extended package on the code to see if this would fix the problem, but none corrected the problem, or would result in other runtime errors. In short, incorrect package usage is often a fatal problem in the generated code [24, 12].
  • Superficial understanding of APIs. This is best illustrated by the Dishes MS using https://api-ninjas.com/api/nutrition. The spec purposely changes the names of the nutritional elements used in the MS from those provided by the API. Furthermore, for some nutritional elements, such as protein, the API returns the string “Only available for premium subscribers”. These elements are not required by the app, but the code must know that not every element in the array will be numeric or it throws an exception. This level of API understanding was beyond the LLM. To overcome this problem, we developed a new prompt that explicitly states the details of the API. Although using this prompt the generated code was able to address these API complexities, the test results remained largely the same due to other issues, like those enumerated below. Finding appropriate APIs and correctly implementing them is another recognized LLM issue [14, 16, 22, 25].
  • Lacking or incorrect datatype conversions. For instance, the Dishes implementation checks that the type of a parameter is a list. But it first needs to use json.loads to convert the string data to JSON. Its failure to do so results in a runtime error and the POST request fails. This has a cascading effect - since the POST fails, subsequent GET requests fail. This error is pervasive and contributes to the very low passing rate for some of the services.
    A similar issue is converting between REST APIs parameters, passed as a string, and the type required to use in an API invocation. For example, a resource (e.g., a specific dish) is identified by its string ID in the REST API. But MongoDB uses, by default, ObjectIDs to identify Mongo documents. Failure to convert between these datatypes was a common error in the library application using GPT-3.5. GPT-o3-mini largely understands the need for this conversion and avoids these mistakes. However there are still times when it fails to perform the conversion due to subtleties of the Mongo API. Specifically, the generated code: dishes_collection.insert_one(dish_record) followed by jsonify(dish_record) causes a run-time error:
    TypeError: Object of type ObjectId is not JSON serializable. This is because by default Mongo adds an ObjectId to the dish_record. One needs to remove this added field or convert it to a string before applying the jsonify() method. Mistaken LLM type inference is a problem for LLMs [19].

### v2

@@ 5.2 Installing packages
The Docker Test Agent runs the generated code in a Docker container. The Dockerfile for that container contains a command to install dependencies in the requirements.txt file using pip. The agent generates the requirements.txt file by scanning the generated code for import statements [^5] In our prompt we asked the LLM to document which packages to install. But we found that it would make mistakes and it was a better solution to install packages based upon the import statements. . This was problematic as pip would fail when installing a package that comes with the standard python distribution (such as uuid or json). To get around this problem, we omitted packages included in the Python standard library.
In the restaurant application (Section 5.4), the generated code for the authentication MS would always import the jwt package, but it required the installation of pyJWT. (To its credit, the LLM would note this in a comment.) For this and other similar issues, our code substitutes the package name in the import statement with the correct package. Nonetheless, as we show below and documented by other research [28, 14, 12], choosing the right package and version is still problematic.
@@ 5.3 The Public Library Application
Our first MSBA is a public library application comprised of 4 MSs: Cardholders, Books, Borrows, and Logs. The Cardholders service was given in Figure 1. Books is responsible for managing the books in the library and uses the Google Books API service (see https://developers.google.com/books/docs/v1/using) to supply additional information about a book. Borrows is responsible for tracking books on loan, and Logs keeps a record of each book that is borrowed from the library. The specification of these MSes can be found on-line (see footnote 2).
We test this application using 3 different LLMs to power the Code and Regen Agents: GPT-3.5-16K, GPT-4o-mini, and GPT-o3-mini. For each LLM we performed two sets of experiments, one using a 0-shot prompt, and one using a $1$-shot prompt [6]. The $1$-shot prompt, given in Appendix C, provides example code using Flask and MongoDB for a small MS. For each LLM and each MS of the MSBA we repeat each experiment 15 times. For GPT-3.5 and GPT-4o-mini we perform 5 experiments for each temperature $0.0$, $0.3$, and $0.5$. GPT-o3-mini did not support different temperature settings and hence all experiments used the default temperature. The Docker Test Agent tests the MSBA with unit tests for each MS. There are 15 unit tests for Cardholders, 14 for Books and Borrows, and 7 for Logs. The latter has fewer tests because Logs only allows external actors to perform GET requests; they cannot create, delete or modify logs on their own. For each experiment, if the LLM generated code does not pass all the unit tests, the Docker Regen Agent prompts the LLM to reflect on the nature of the errors and then prompts it to regenerate code, as explained above. We only do this once (MaxGen = $2$ in Figure 2). We refer to the initial generated code as V0 and the regenerated code as V1.
@@ 5.3.1 Results
Tables 1 and 2 shows the results of our experiments on GPT-3.5 with the 0-shot and $1$-shot prompts respectively. The $1^{st}$ column is the name of the MS. As not all generated code was executable, the $2^{nd}$ column lists the fraction of the 15 experiments that generated executable V0 code permitting execution of the unit tests. The $3^{rd}$ column lists the number of experiments that generated executable V1 code. The $4^{th}$ and $5^{th}$ columns present the percent of unit tests that passed using V0 and V1 code respectively, averaged over all experiments that produced executable code [^6] We do not consider unexecutable generated code in this percentage. We wanted our results to focus on how often executable code was functionally correct. . The $6^{th}$ column is of the form $n_{1}$-$n_{2}/t$, where $n_{1}$ ($n_{2}$) is the highest number of unit tests passed in any experiment for V0 (V1) and $t$ is the total number of unit tests for that MS. The last column is the percent of experiments in which the number of unit tests passed by V1 was greater than the number of unit tests passed by V0, considering only those experiments in which V1 was executed.
Name # V0 # V1 % V0 % V1 Highest % V1
@@
Table 1: Public library MSBA results using GPT-3.5 with 0-shot prompt
Name # V0 # V1 % V0 % V1 Highest % V1
@@
Table 2: Public library MSBA results using GPT-3.5 with 1-shot prompt
@@
Table 3: Public library MSBA results using GPT-o3-mini with 0-shot prompt
@@
Table 4: Public library MSBA results using GPT-o3-mini with 1-shot prompt
We omit the results for GPT-4o-mini due to space considerations. In general it performs better than GPT-3.5 but not as good as GPT-o3-mini. See Appendix E.
@@ 5.3.2 Discussion
A third observation, based upon column 4 of Tables 2 and 4, is that providing an example (1-shot prompt) significantly improves performance of GPT-3.5 but mostly degrades performance with GPT-o3-mini. This may be the case with reasoning models like GPT-o3-mini [^7] For instance, the creators of DeepSeek -R1 report that “When evaluating DeepSeek-R1, we observe that it is sensitive to prompts. Few-shot prompting consistently degrades its performance. Therefore, we recommend users directly describe the problem and specify the output format using a zero-shot setting for optimal results.” [9] See also [31]. , and especially in our case, where performance is high without the example, and the example can cause overfitting.
@@
A general comment is warranted on these statistics; namely, not all unit tests are equal. Some unit tests may require complicated logic while others check very simple REST patterns that do not even require knowledge of the MS. For instance, one of the unit tests checks that a POST request with a non-JSON payload fails. Hence while percentage of unit tests passed still gives us an overall idea on the strength of the LLM, we require a deeper analysis examining which tests pass or fail to obtain additional insight (see Section 5.7).
With respect to temperature, while the $0.3$ setting often performs best for V0 and $0.5$ for V1, the differences in performance between different temperatures is usually less than $10\%$. See Appendix E for details.
@@ 5.4 The Restaurant Application
Given the poor results of GPT-3.5 on the Library app, we use only GPT-4o-mini and GPT-o3-mini for our experiments on this MSBA. For similar reasons, we only experiment with a 0-shot prompt. For GPT-4o-mini we performed 20 experiments for each MS. For the ratings MS, 8 of the experiments were at temperature $0.0$, 7 were at temperature $0.3$ and 5 were at temperature $0.5$. For all other microservices, there were 10 experiments at temperature $0.0$, and 5 experiments at each of the temperatures $0.3$ and $0.5$. For GPT-o3-mini we performed 15 experiments for each MS. As mentioned above GPT-o3-mini did not support different temperatures so we used the default temperature.
Each experiment consists of 9, 15, 10 and 8 unit tests for the Authentication, Dishes, Profile, and Ratings MSes respectively. As above, for each experiment, if the LLM generated code does not pass all unit tests, the Regen Agent reflects on the nature of the errors and then regenerates code.
@@ 5.6 Description Complexity
Before analyzing the results of these experiments, we first state the description complexity of the restaurant and library MSBAs, given in Table 7. The description size is for the 0-shot prompt. The 1-shot prompt adds $490$ words.
The Restaurant MSBA has a description size about .36.36 larger than the Library application, and has twice the number of dependencies, although the package complexity is the same. The resource complexity of the Restaurant application is slightly larger. This at least partially explains the difference in LLM performance between the library and restaurant MSBAs.
Name Description # Dependencies Avg # Packages Primitive-Coll LLM assigned
size (words) per MS Resources difficulty
Library 1399 3 4 22-4 3
Restaurant 1905 6 4 25-5 4
@@
The LLM-as-a-Judge score, provided by Claude Sonnet 4, confirms this conclusion. It assigns the library application a difficulty of 3 (out of 5) while the restaurant application received a complexity score of 4. It says that the library app has “moderate complexity” and that it is not overly complex because of the “Clear, well-defined requirements”, “Standard CRUD operations”, “Common technology stack”, “Simple data models”, and “Limited business logic”. Factors that add complexity include “Inter-service communication”,“External API integration”, “Business rule enforcement”, “Date calculations”, “Error handling”, and “Distributed data consistency”.
For the restaurant MSBA, the LLM Judge provides the following “High Complexity Elements” contributing to the 4 rating: “Multi-service Architecture”, “Authentication & Authorization”, “External API Integration”, “Complex Business Logic”, “Database Operations”, “Error Handling”, and “Query Processing”. It did not give it a 5 rating because of “Well-defined specifications with clear requirements”, “Standard technologies (Python, MongoDB, REST APIs)”, “No real-time processing or complex distributed systems patterns” and “No advanced security beyond basic token auth”. In Appendix B we provide the LLM-as-a-Judge prompt and the comments provided by the LLM to explain its score for the restaurant application. The comments for the library app can be found on-line (see footnote 2).
@@ 5.7 Discussion
From Table 5 we see that GPT-4o-mini does extremely poorly on the restaurant application, and while GPT-o3-mini (Table 6) does better, it still never performs above $25\%$ passing unit tests with V0 code except for the Authentication MS, where it achieves $65.8\%$. We also note that neither V0 nor V1 code ever achieves a perfect score, with the exception of the Authentication service. In 6 out of the 11 V1 testable implementations, it passed all $9$ unit tests.
To get a deeper understanding of the cause of the errors, we examined the run-time error-logs, the GPT-o3-mini reflection results, as well as the generated code. Among the persistent problems are the following:
  • Incorrect package usage. For instance, even though we installed pyJWT (see Section 5.2 ), the authentication code would often use this package incorrectly. Along similar lines, Ratings would consistently import the flask_jwt_extended package. A typical error message was the following:
@@
    We tried multiple versions of the flask_jwt_extended package but none corrected the problem, or would result in other runtime errors. In short, incorrect package usage is often a fatal problem in the generated code [28, 14].
  • Superficial understanding of APIs. This is best illustrated by the Dishes MS using https://api-ninjas.com/api/nutrition. The spec purposely changes the names of the nutritional elements used in the MS from those provided by the API. Furthermore, for some nutritional elements, such as protein, the API returns the string “Only available for premium subscribers”. These elements are not required by the app, but the code must know that not every element in the array will be numeric or it throws an exception. This level of API understanding was beyond the LLM. To investigate further, we developed a new prompt that explicitly states the details of the API. Although using this prompt the generated code was able to address these API complexities, the test results remained largely the same due to other issues, like those enumerated below. Finding appropriate APIs and correctly implementing them is another recognized LLM issue [18, 20, 26, 29].
  • Lacking or incorrect datatype conversions. Surprisingly, the generated Dishes implementation has difficulty converting the JSON payload to a string. Its failure to do so results in a runtime error and the POST request fails. This has a cascading effect - since the POST fails, subsequent GET requests fail. This error is pervasive and contributes to the low passing rate for some of the services.
    A similar issue is converting between REST APIs parameters, passed as a string, and the type required to use in an API invocation. For example, a resource (e.g., a specific dish) is identified by its string ID in the REST API. But MongoDB uses, by default, ObjectIDs to identify Mongo documents. Failure to convert between these datatypes was a common error in the library application using GPT-3.5. GPT-o3-mini generally performs this conversion. However there are still times when it fails to do so due to subtleties of the Mongo API. Specifically, the generated code: dishes_collection.insert_one(dish_record) followed by jsonify(dish_record) causes a run-time error:
    TypeError: Object of type ObjectId is not JSON serializable. This is because by default Mongo adds an ObjectId to the dish_record. One needs to remove this added field or convert it to a string before applying the jsonify() method. Type inference is a problem for LLMs [22].
@@
6 Fine-grained experiments
We conjecture that using a fine-grained approach to generate complex code can improve the results. Specifically, instead of generating code for the entire service, the Code and Regen Agents generate code for each request type (e.g., POST, GET, …) independently. They then merge the generated code into the GT code for that service. For example, if the Code Agent generates code for the POST request, it will then replace the POST request code in the GT file with the generated code but the rest of the GT code remains the same. The Docker Test Agent still tests the entire application, but the results will be solely dependent on the generated POST code. The value of this approach is that it:
  • Adopts a divide-and-conquer approach to code generation, allowing the LLM to focus on generating a smaller piece of code for each prompt.
  • Facilitates further understanding of the cause of the errors in the generated code, as the generated code is only for a specific request.
  • Removes the possibility that an error in one request type (e.g., GET) is due to an error in a previous request of another type (e.g., POST).
The GT code contains assumptions that all services must share, for instance, the name of the Mongo collection that stores service resources. Furthermore, there are cases where merging the generated code for a specific request into the existing GT implementation can cause conflicts if care is not taken. For this reason the prompts used by the Code and Regen agents include guidance on how to generate the code: it specifies the name of the Mongo collection, it requests that a wrapper be used for authentication, and it declares the name of packages that are available to import. Appendix D provides the Code Agent prompt for generating POST request code for the dishes service. The prompts for other requests types and other services are similar.
The results of the fine-grained experiments for the dishes and ratings services of the restaurant MSBA are given in figures 8 and 9 respectively. All experiments were performed using GPT-o3-mini. In these tables each row except for the last one presents the results for generated code of the request type given in the $1^{st}$ column. The $2^{nd}$ column lists the number of experiments performed. The $3^{rd}$ column lists the number of unit tests that explicitly test that request type in each experiment [^8] For example, POST requests that explicitly test the POST function are included in these numbers. POST requests that are used to set up a test on another request type, such as DELETE, are not included in these numbers. and the $4^{th}$ column and $5^{th}$ columns gives the number and percent of those tests that passed for V0 code. The last two columns state the number and percent of those tests that passed for V1 code. The last row presents the average number of passing unit tests for V0 and V1, averaged over the rows above it. For Table 9, we consider it as if $100$ is written in the last column, since $100\%$ of the V0 tests passed and therefore no V1 code was generated.
We first compare the fine-grained results on the dishes service to the coarse-grained results for that service. We note that the latter results in $23.6\%$ and $25.8\%$ for V0 and V1 code respectively (Table 6, row 2):
  • The V0 results for the fine-grained generation of POST requests for the dishes service (Table 8, row 1, col 5) yields approximately the same percent of V0 passing tests as for the dishes service in the coarse grained experiments.
  • The V0 results for the fine-grained generation of the GET and DELETE requests (rows 2 and 4) show remarkable improvement in passing percent over the coarse-grained experiments. PUT (row 3) is lower, but each experiment has a single PUT test and therefore one cannot overly rely on its results, given the scarcity of data.
  • All the requests, including POST, show significantly better results on V1 code than the coarse-grained results.
  • The previous observations leads one to infer that (1) POST requests are harder to generate, and because the POST code often fails in the coarse-grained service generation, the subsequent requests will likewise fail. (2) Generating smaller pieces of code (at the request level) pays off. The average passing percent across all 4 request types for V0 and V1 code is $48.5\%$ and $77.8\%$ respectively (Table 8, last row). The is significantly better than $23.6\%$ and $25.8\%$ for coarse-grained generation. The improvement between V0 and V1 percents is also very much higher. However, these observations are tempered by the knowledge that the fine-grained code is easier for the LLM to generate as we supply the LLM with information that is not given in the coarse-grained prompts (authentication code, import statements, and Mongo collection name.).
  • A good example fine-grained generation improvement over coarse-grained generation can be seen by the issue we cited in Section 5.7 “Lacking or incorrect datatype conversions”. We noted there that the generated code dishes_collection.insert_one(dish_record) did not remove the Mongo _id before the jsonify call. This was true for both the V0 and V1 code in the coarse-grained experiments. While this was true for the V0 code in fine-grained experiments, for V1 code, $10/12$ times it corrected this error.
  • Generating fine-grained requests for the ratings service is less beneficial than for the dishes service. Specifically, generating the POST V0 request code fails in all experiments, as seen in Table 9 row 1. Surprisingly, generating GET requests (row 2) likewise proves hard for the LLM in V0 code. However, V0 PUT code is correct for every experiment (row 3).
  • Nonetheless, there is still significant improvement in using fine-grained instead of coarse-grained generation for the ratings service. The average passing percent across all 3 request types for V0 and V1 code is $33.3\%$ and $63.2\%$ respectively (Table 9, last row). The is much better than $10.8\%$ and $39.2\%$ percent for coarse-grained generation (Table 6, row $4$, columns $4$ and $5$).
Request # experiments # tests # V0 % V0 # V1 % V1
passed passed passed passed
POST 12 7 20/84 23.8 38/84 45.2
GET 13 5 51/65 78.5 58/65 89.2
PUT 13 1 0/13 0 10/13 76.9
DELETE 12 2 22/24 91.7 4/4 100
Overall Avg 12 48.5 77.8
Table 8: Results for fine-grained generated code for the dishes restaurant service
Request # experiments # tests # V0 % V0 # V1 % V1
passed passed passed passed
POST 12 4 0/48 0 11/48 22.9
GET 12 3 0/36 0 24/36 66.7
PUT 12 1 12/12 100 - -
Overall Avg 33.3 63.2
Table 9: Results for fine-grained generated code for the ratings restaurant service
In summary, the results indicate that fine-grained request generation is a promising direction to pursue. It also shows that additional focus needs to be given to generating the POST request as it is the most difficult request type to generate, and if the POST code does not work properly, the other requests will also fail.